OpenStudy (anonymous):

is 17.142857142857142 the correct anwser for Find each percent of change and round to the nearest tenth. 70 to 82

OpenStudy (anonymous):

change is 82 - 70 = 12 Percent change is 12/70 = ?

OpenStudy (anonymous):

is 17.142857142857142 the correct anwswer

OpenStudy (anonymous):

12/70 = 0.171428571 ... nice job!

OpenStudy (anonymous):

thanks metal rewared and fanned